A five-digit number is formed by arranging 2, 3, 4, 5, and 6 randomly without repetition. Find the
probabilities of the following events,
The number is an even number.

Answer by jitendra_maths(6)

given digits are 2,3,4,5,6
to make an even number, unit place digit must be filled by either 2, 4 or 6
total numbers formed by keeping 2 at unit place = 4! = 4x3x2x1 = 24
similarly numbers formed by keeping 4 at unit place = 24
and numbers formed by keeping 6 at unit place = 24
total even numbers = 24=24+24 = 72
total number of ways by arranging the digits = 5! = 5x4x3x2x1 = 120
Henc, Probability of an even number = 72/120 = 3/5